# VisuRiddles: Fine-grained Perception is a Primary Bottleneck for Multimodal Large Language Models in Abstract Visual Reasoning

### v1

@@ 4 Experiment
In this section, we present extensive experiments on advanced MLLMs on the challenging VisuRiddles benchmark. We first describe the experimental setup for both training and evaluation in Sec. 4.1. Subsequently, we demonstrate a comprehensive evaluation of MLLMs in Sec. 4.2. We further include visualization examples that illustrate model behaviors and failure modes in Sec. 4.3. Finally, we conduct ablation studies to validate the efficacy of our synthesis engine in Sec. 4.4.
@@ 4.1 Experimental Setup
Training details. We fine-tune the open-source model Qwen2.5-VL-7B using 80k perceptually annotated samples generated by our Riddles Synthesis Framework. Training is performed with a learning rate of 5e-6, using the Adam optimizer, a batch size of 16, and 8 NVIDIA A800 GPUs.
Evaluation Settings. We evaluate seventeen MLLMs, including eleven open-source and six advanced proprietary models. The open-source group consists of MiniCPM-V 2.6 [7], DeepSeek-VL2 [67], Qwen2.5-VL [68], InternVL2.5 [5] and InternVL2.5-MPO (Mixed Preference Optimization) [69], which is designed to improve reasoning capabilities via reinforcement learning-based optimization. The proprietary group includes GPT-4o (20241120) [70], o1 [71], Claude-3-7-Sonnet (20250219) [72], Claude-3-7-Sonnet-Thinking (20250219) [72], Gemini-2.0-Flash-Thinking (20250121) [73], and Gemini-2.5-Pro (20250325) [74]. Notably, GPT-o1, Claude-3-7-Sonnet-Thinking, and Gemini-2.5-Pro are equipped with specialized “thinking” modes aimed at enhancing multi-step reasoning. To further explore reasoning capabilities, CoT prompting is applied to GPT-4o, Qwen2.5-VL-72B, and InternVL2.5-78B-MPO, selected for their reasoning capacity, model scale, or lack of built-in “thinking” mode. Inference for all models is performed using the vLLM framework with unified decoding configurations. Human reference answers are obtained from the Fenbi Educational Platform, a widely used source for civil service exam preparation.
Table 3: Evaluation of various models on VisuRiddles. The superscript number indicates the number of answer options (e.g., “4” denotes a 4-choice question where one option is correct; $\infty$” represents unlimited possible options). The best results are marked bold and the second results are underlined.
[…]
Human - 61.3 60.9 67.5 67.9 58.8 - - 61.9 63.0
@@
Minicpm-V-2.6 8B 22.2 23.1 38.9 23.1 27.4 1.0 11.0 28.2 21.9
InternVL2.5-8B 8B 23.0 29.9 27.8 31.7 28.0 0.0 8.0 31.0 22.7
InternVL2.5-8B-MPO 8B 25.7 29.9 20.0 27.9 27.4 0.0 10.0 31.0 22.4
Deepseekvl2 27B 28.4 29.1 22.2 22.1 24.2 0.0 8.0 18.3 21.0
InternVL2.5-38B 38B 22.6 26.5 28.9 27.9 26.8 1.0 9.0 31.0 21.9
InternVL2.5-38B-MPO 38B 24.5 23.1 33.3 29.8 27.4 0.0 8.0 25.4 22.1
Qwen2.5VL-32B 32B 25.7 23.9 24.4 29.8 28.7 1.0 9.0 25.4 22.1
InternVL2.5-78B 78B 21.8 31.6 30.0 27.9 26.8 0.0 15.0 22.5 22.3
InternVL2.5-78B-MPO 78B 26.8 31.6 25.6 26.9 24.8 0.0 13.0 23.9 22.7
InternVL2.5-78B-MPO(cot) 78B 26.1 24.8 27.8 33.7 21.7 0.0 16.0 21.1 22.2
Qwen2.5VL-72B 72B 24.5 23.1 23.3 29.8 22.3 2.0 12.0 29.6 21.3
Qwen2.5VL-72B(cot) 72B 26.4 25.6 24.4 32.7 26.8 0.0 18.0 31.0 23.7
@@
GPT-4o - 24.5 22.2 28.9 25.0 22.3 0.0 17.0 25.4 21.2
GPT-4o(cot) - 28.7 28.2 34.4 27.9 22.9 0.0 11.0 33.8 23.9
GPT-o1 - 31.0 28.2 33.3 25.0 26.1 0.0 15.0 26.8 24.5
claude-3-7-sonnet - 25.3 23.1 30.0 26.0 25.5 1.0 12.0 28.2 22.0
claude-3-7-sonnet-thinking - 21.1 26.5 41.1 30.8 26.8 9.0 13.0 26.8 23.8
Gemini-2.0-flash-thinking - 17.6 12.8 35.6 26.9 17.2 10.0 11.0 25.4 18.7
Gemini2.5-pro - 26.4 32.5 44.4 29.8 23.6 15.0 14.0 33.8 26.8
@@
Baseline(Qwen2.5VL-7B) 7B 23.4 28.2 28.9 23.1 21.7 0.0 14.0 23.9 20.9
PRS-Qwen2.5VL-7B 7B 27.2 35.7 41.4 38.3 30.3 60.0 95.0 21.3 43.7
@@ 4.2 Main Results
Performance of Open-Source MLLMs. Most open-source MLLMs demonstrate limited performance on VisuRiddles, with accuracies across core categories often close to random choice and consistently failing on the two high-level reasoning tasks, which reveals a substantial gap between current model capabilities and human performance in AVR.
Performance of Commercial Models with Thinking Mode. Several commercial MLLMs with ‘thinking’ mode, such as GPT-o1, Claude-3-7-Sonnet-Thinking, and Gemini2.5-Pro, achieve comparatively improved performance on VisuRiddles. Their advantage is particularly notable in structured tasks like Sudoku and Raven. These results indicate that AVR can benefit from advanced reasoning capabilities. However, the overall performance remains unsatisfactory, suggesting that reasoning enhancements alone are insufficient to address the core challenges of this task.
Effectiveness of Model Scaling and CoT Prompting. Strategies such as increasing model size and applying CoT prompting, though commonly used to enhance reasoning, are ineffective for AVR. Models with larger parameter scales do not consistently outperform smaller ones, and CoT yields only limited improvements on select tasks. These findings indicate that scaling in model parameter or inference time is insufficient to address the core challenges of AVR.
Effectiveness of Our Perception-Enhanced Model. Our perception-enhanced model significantly outperforms other MLLMs across all categories of VisuRiddles, particularly in high-level reasoning tasks, Sudoku reasoning and Raven reasoning, demonstrating the model’s ability to handle structurally complex visual patterns. Importantly, these gains are achieved without changing model architecture or size, but solely through training on data with perceptual description, highlighting that the improvement arises from enhanced perceptual abstraction rather than stronger reasoning capacity. These results affirm the critical role of fine-grained perceptual supervision in abstract visual reasoning and emphasize the need to model intermediate perceptual processes for better performance on structurally complex tasks.
Figure 4: Case study comparing different reasoning strategies on a VisuRiddles example. (a) QVQ-72B reflects a flawed loop under the ”thinking” mode. (b) Qwen2.5-VL-72B with CoT prompting exhibits incorrect perceptual understanding. (c) Our perception-enhanced model generates structured reasoning grounded in accurate perceptual abstraction, yielding the correct answer.
@@ 4.3 Visualization Results
To further analyze the reason that MLLMs underperform on AVR task, we examine two widely adopted reasoning enhancement strategies: inference-time scaling and CoT prompting. We conduct experiments on QVQ-72B-Preview (equipped with a ”think” mode), Qwen2.5-VL-72B (with CoT prompting), and our perception-enhanced model Qwen2.5-8B.
Fig. 4 visualizes a representative case of a broadly observed phenomenon uncovered during our experiments. The results reveal clear differences in the reasoning behaviors across these models. The inference-time scaling approach, which leverages ”think”-style prompting without any perceptual grounding, frequently leads to verbose yet logically inconsistent reasoning, often ending in incorrect answers. The CoT prompting, while occasionally yielding the correct answer, fails to guide the model along a valid reasoning path due to inaccurate visual perception. In such cases, correct predictions can be largely attributed to the single-choice format rather than abstract visual understanding. This observation aligns with CoT-based models perform close to random guessing on basic inference tasks, but struggle significantly on high-level inference tasks, shown in Tab. 3. In contrast, our approach leverages fine-grained perception to produce concise, visually grounded reasoning and accurate answers.
These results underscore the importance of perception in abstract visual reasoning. Inference-time scaling and CoT prompting alone cannot compensate for the lack of perceptual grounding, which often leads to incorrect or inconsistent reasoning. In contrast, introducing fine-grained perceptual supervision enables the model to build accurate and visually aligned reasoning paths, supporting reliable performance on complex visual tasks.
@@ 4.4 Ablation Study
Ablation Study on Bottlenecks in Abstract Visual Reasoning. To investigate whether the performance gap in abstract visual reasoning stems from perceptual limitations rather than reasoning ability, we conduct an ablation study using our perception-annotated synthetic dataset. We evaluate two representative MLLMs, GPT-4o (closed-source) and Qwen2.5VL-72B (open-source), on a subset of our synthetic dataset under two input settings: (i) raw abstract graphics (V), and (ii) structured perceptual descriptions (P).
The experimental results in Tab. 4 show that both models perform poorly when given only raw visual inputs, across both the five core perceptual dimensions and the two high-level reasoning tasks. Taking Sudoku reasoning as an example, this task requires recognizing digit values and their positions within a grid, and inferring a globally consistent solution. Under the Visual setting, both GPT-4o and Qwen2.5VL-72B fail to handle such constraint-based reasoning, while in the perception setting, accuracy increases from near-zero to 15% and 65%, respectively. This improvement primarily stems from structured perceptual descriptions compensating for the MLLMs’ limited perceptual abstraction capabilities. The difficulty stems from the nature of Sudoku images: small digits, dense layout, and absent semantic cues make them hard for MLLMs to interpret, though humans process such structure effortlessly through rapid visual scanning.
Table 4: Performance Impact of Perceptual Descriptions vs. Visual Inputs in MLLMs. Where ”V” denotes inputting abstract graphics directly, while “P” refers to replacing graphics with structured perceptual descriptions.
Model Num Styl Attr Posit Spat Sudo Rav Avg
@@
Table 5: Ablation study on dataset setting
Model Num Styl Attr Posit Spat Sudo Rav Other Avg
Baseline 23.4 28.2 28.9 23.1 21.7 0.0 14.0 23.9 20.9
+20K 24.5 (+1.1) 29.8 (+1.6) 31.2 (+2.3) 26.1 (+3.0) 23.8 (+2.1) 19.0 (+19.0) 40.0 (+26.0) 22.0 (-1.9) 26.4 (+5.5)
+40K 25.9 (+2.5) 31.8 (+3.6) 34.7 (+5.8) 29.7 (+6.6) 26.4 (+4.7) 30.0 (+30.0) 66.0 (+52.0) 23.5 (-0.4) 30.9 (+10.0)
+60K 26.7 (+3.3) 33.8 (+5.6) 37.8 (+8.9) 33.5 (+10.4) 28.6 (+6.9) 53.0 (+53.0) 87.0 (+73.0) 24.8 (+0.9) 37.2 (+16.3)
Full data 27.2 (+3.8) 35.7 (+7.5) 41.4 (+12.5) 38.3 (+15.2) 30.3 (+8.6) 60.0 (+60.0) 95.0 (+81.0) 21.3 (-2.6) 43.7 (+22.8)
Ablation Study on Perceptually Annotated Synthetic Data. To evaluate the effectiveness of our synthesized data in enhancing model performance on abstract visual reasoning, we conduct a series of ablation studies with incremental training data. The results are shown in Tab. 5. Increasing the amount of synthetic data with perceptual annotations leads to consistent improvements across all dimensions except for the ”Other” category, demonstrating that perceptually structured supervision provides effective guidance for internalizing abstract visual patterns that are difficult to learn from isolated answer-based supervision alone. This further highlights the importance of modeling intermediate perceptual processes in abstract visual reasoning.
@@ 5 Conclusion
In this work, we introduce VisuRiddles, a benchmark that systematically evaluates MLLMs’ AVR abilities across diverse, structurally rich tasks, along with an automated data synthesis framework that generates controllable, perceptually annotated reasoning problems. Experimental results demonstrate that even state-of-the-art MLLMs perform only marginally above random chance on VisuRiddles, exhibiting a substantial gap compared to human performance. Our analysis reveals that their limitations stem from inadequate perceptual abstraction rather than reasoning capacity, and that training with our synthetic data improves performance and interpretability. In sum, VisuRiddles enables fine-grained evaluation of reasoning, and our framework offers scalable data support for future training. More importantly, these findings underscore a central insight: enhancing perceptual abstraction is critical to advancing abstract visual reasoning in MLLMs.

### v3

@@ 5.1 Experimental Setup
We use Qwen2.5-VL-7B [4] as the baseline model to develop PAVR. In the supervised fine-tuning stage, the model is trained for 20 epochs using 20K synthesized AVR instances generated by the VisuRiddles Synthesizer. Subsequently, in the GRPO stage, optimization is performed using 4K synthesized instances over 40 epochs. All experiments are conducted on a setup comprising 8 NVIDIA A800 80G GPUs. Evaluation settings are summarized in Appendix C.
Table 2: Evaluation result on VisuRiddles benchmark. The superscript number indicates the number of answer options (e.g., “4” means a 4-choice question with one correct option; * represents a vast combinatorial solution space). The best results are marked bold and the second results are underlined.
Model Param Num4 Styl4 Attr4 Posit4 Spat4 Sudo* Rav8 Other4 Avg
Human - 61.3 60.9 67.5 67.9 58.8 - - 61.9 -
@@
Minicpm-V-2.6 8B 21.6 26.5 23.7 31.5 23.1 0.0 10.0 24.6 20.6
InternVL2.5-8B 8B 21.2 13.7 30.9 23.4 23.1 0.0 0.0 29.0 18.1
InternVL2.5-8B-MPO 8B 22.0 23.9 27.8 25.2 21.8 0.0 3.0 27.5 19.4
Deepseekvl2 27B 21.2 27.4 18.6 18.0 17.3 0.0 15.0 13.0 17.4
InternVL2.5-38B 38B 22.0 19.7 29.9 26.1 29.5 1.0 19.0 30.4 22.3
InternVL2.5-38B-MPO 38B 26.0 20.5 19.6 27.0 27.6 0.0 18.0 31.9 22.1
Qwen2.5VL-32B 32B 23.2 20.5 30.9 27.0 29.5 0.0 40.0 24.6 24.5
InternVL2.5-78B 78B 26.0 25.6 27.8 29.7 23.7 0.0 16.0 27.5 22.7
InternVL2.5-78B-MPO 78B 25.6 30.8 24.7 27.9 26.3 0.0 12.0 20.3 22.2
InternVL2.5-78B-MPO(cot) 78B 23.2 23.1 24.7 28.8 22.4 0.0 11.0 23.2 20.3
Qwen2.5VL-72B 72B 23.6 23.1 19.6 30.2 26.9 0.0 62.0 23.9 25.9
Qwen2.5VL-72B(cot) 72B 27.2 26.5 24.7 19.8 28.2 0.0 55.0 23.2 26.0
Qwen3-VL-235B-A22B-Instruct 235B 31.2 28.2 36.1 34.2 26.3 29.0 53.0 27.5 32.6
Qwen3-VL-235B-A22B-Thinking 235B 31.2 29.9 44.3 33.3 30.1 33.0 49.0 39.1 34.9
Dots.vlm1 671B 30.4 31.6 38.1 38.7 35.3 1.0 20.0 33.3 29.2
@@
GPT-4o - 23.2 21.4 29.9 27.9 28.2 0.0 14.0 24.6 21.8
GPT-4o(cot) - 22.4 23.9 32.0 29.7 30.1 0.0 22.0 29.0 23.7
o3 - 28.4 40.2 41.2 27.0 21.8 0.0 25.0 33.3 27.0
claude-3-7-sonnet - 24.0 27.4 27.8 29.7 28.2 4.0 43.0 31.9 26.5
claude-3-7-sonne(cot) - 24.4 36.8 27.8 29.7 25.6 3.0 36.0 27.5 26.2
claude-3-7-sonnet-thinking - 24.0 24.8 33.0 23.4 25.6 3.0 24.0 26.1 23.2
Gemini-2.5-flash-thinking - 21.2 30.8 27.8 21.6 16.7 17.0 16.0 23.2 21.5
Gemini2.5-pro - 31.6 31.6 48.5 26.1 30.1 39.0 30.0 44.9 33.9
GPT-5 - 30.8 30.8 38.1 32.4 30.8 2.0 29.0 31.9 28.7
@@
Baseline(Qwen2.5VL-7B) 7B 24.4 28.2 23.7 22.5 25.0 0.0 48.0 24.6 24.6
PAVR-SFT 7B 31.2 31.6 44.3 31.5 45.5 43.0 61.0 39.1 39.5
PAVR 7B 39.6 39.3 50.5 39.6 51.9 46.0 65.0 55.1 46.8
@@ 5.2 Main Results
Performance of Open-Source MLLMs. Most open-source MLLMs demonstrate limited performance on VisuRiddles, with accuracies across core categories often close to random choice and consistently failing on the two high-level reasoning tasks, which reveals a substantial gap between current model capabilities and human performance in AVR. Notably, the newly released Qwen3-VL-235B-A22B-Instruct and Qwen3-VL-235B-A22B-Thinking exhibit markedly stronger results, though such improvements remain exceptions among open-source models.
Performance of Commercial Models with Thinking Mode. Several commercial MLLMs with ‘thinking’ mode achieve comparatively improved performance on VisuRiddles. Their advantage is particularly notable in structured tasks like Sudoku and Raven. These results indicate that AVR can benefit from advanced reasoning capabilities. However, the overall performance remains limited.
Effectiveness of Model Scaling and CoT Prompting. Strategies such as increasing model size and applying CoT prompting, though commonly used to enhance reasoning, are ineffective for AVR tasks. Models with larger parameter scales do not consistently outperform smaller ones, and CoT yields only limited improvements on select tasks. These findings indicate that scaling in model parameter or inference time is insufficient to address the core challenges of AVR.
Effectiveness of PAVR. PAVR significantly outperforms other MLLMs across all categories of VisuRiddles, especially in high-level reasoning tasks, demonstrating the model’s ability to handle complex visual patterns. Notably, compared to the baseline, the significant improvement achieved by training on data with perceptual descriptions through SFT highlights the critical role of fine-grained perceptual ability in AVR tasks. On the other hand, the application of RL further enhances performance across various dimensions on VisuRiddles, with the improvements primarily stemming from enhanced reasoning capabilities through strategic guidance.
We also provide the results of PAVR on the VisuLogic [80] benchmark in  D. The experimental results further support our conclusions.
@@ 5.3 Visualization Results
Figure 5: Case study comparing different reasoning strategies on a VisuRiddles example. (a) QVQ-72B reflects a flawed loop under the "thinking" mode. (b) Qwen2.5-VL-72B with CoT prompting exhibits incorrect perceptual understanding. (c) PAVR exhibits accurate perception and coherent reasoning, ultimately arriving at the correct answer.
To further analyze the reason that MLLMs underperform on AVR task, we examine two widely adopted reasoning enhancement strategies: inference-time scaling and CoT prompting. We conduct experiments on QVQ-72B-Preview (equipped with a "think" mode), Qwen2.5-VL-72B (with CoT prompting), and PAVR. Fig. 5 visualizes a representative case of a broadly observed phenomenon uncovered during our experiments. The results reveal clear differences in the reasoning behaviors across these models. The inference-time scaling approach, which leverages "think"-style prompting without any perceptual grounding, frequently leads to verbose yet logically inconsistent reasoning, often ending in incorrect answers. The CoT prompting strategy often suffers from perceptual errors, which cascade into flawed reasoning and ultimately lead to incorrect answers. In contrast, by grounding reasoning in fine-grained perceptual capability, PAVR correctly interprets the riddle and identifies the underlying pattern, ultimately producing the correct answer. Notably, while PAVR occasionally exhibits erroneous perception, we observe a ‘rethink’ phenomenon, where the model re-examines and corrects them, realigning the reasoning trajectory to reach the correct solution.
These results underscore the importance of perception in AVR. Inference-time scaling and CoT prompting alone cannot compensate for the lack of perceptual grounding, which often leads to incorrect or inconsistent reasoning. Conversely, the incorporation of fine-grained perceptual capability enables the model to capture subtle visual cues and sustain reliable performance on AVR tasks.
@@ 5.4 Ablation Study
Ablation Study on Bottlenecks in AVR. To investigate whether the performance gap in AVR stems from perceptual limitations rather than reasoning ability, we conduct an ablation study using our perception-annotated synthetic data. We evaluate two representative MLLMs, GPT-4o (closed-source) and Qwen2.5VL-72B (open-source), on a subset of our synthetic dataset under two input settings: (i) raw abstract graphics (V), and (ii) structured perceptual descriptions (P).
The experimental results in Tab. 3 show that both models perform poorly when given only raw visual inputs. Taking Sudoku reasoning as an example, this task requires recognizing digit values and their positions within a grid, and inferring a globally consistent solution. Under the Visual setting, both GPT-4o and Qwen2.5VL-72B fail to handle such constraint-based reasoning, while in the perceptual descriptions setting, accuracy increases from near-zero to 15% and 65%, respectively. This improvement primarily stems from structured perceptual descriptions compensating for the MLLMs’ limited perceptual abstraction capabilities. The difficulty stems from the nature of Sudoku images: small digits, dense layout, and absent semantic cues make them hard for MLLMs to interpret, though humans process such structure effortlessly through rapid visual scanning.
Table 3: Impact of Descriptions vs. Visual Inputs in MLLMs. Where "V" denotes inputting abstract graphics directly, while “P” refers to replacing graphics with structured perceptual descriptions.
Model Num. Styl. Attr. Posit. Spat. Sudo. Rav. Avg
@@
Ablation Study on Reasoning Enhancements in AVR. To examine the contribution of reasoning-oriented enhancements, we conduct an ablation study on three components: Caption, CoT, and GRPO, as shown in Table 4. We can draw the following conclusions form the experiments: (i) perception serves as the essential foundation, as reasoning enhancements alone contribute only marginal improvements; (ii) relying solely on structured perceptual descriptions for SFT yields limited gains and weak generalization, whereas augmenting SFT with CoT annotations generated by MLLMs alleviates this issue, endowing the model with fine-grained perceptual ability and partial reasoning capability; and (iii) Building on a perception-grounded model, the incorporation of GRPO further unifies perception and reasoning, ultimately delivering substantial performance improvements.
Table 4: Ablation study on Reasoning Enhancements. Where Caption denotes training with fine-grained perceptual descriptions, CoT denotes augmenting training with chain-of-thought reasoning traces that incorporate structured perceptual descriptions generated by MLLMs, and GRPO refers to applying Group Relative Policy Optimization to strength model’s reasoning capability.
Model SFT RL Avg
Baseline (Qwen2.5-VL) 24.6
Baseline + Caption 33.3 (+8.7)
Baseline + GRPO 29.4(+4.8)
Baseline + CoT (PAVR-SFT) 39.5(+14.9)
Baseline + CoT + GRPO (PAVR) 46.8 (+22.2)
@@ 6 Conclusion
In this work, we observe that the primary bottleneck of MLLMs in AVR lies in their overlooked fine-grained perceptual capability. To address this issue, we introduce VisuRiddles, a unified resource that comprises a benchmark for objectively evaluating MLLMs on AVR tasks and a synthesizer for generating training instances with structured perceptual descriptions. Building on this resource, we propose PAVR, which unifies perception and reasoning through supervised fine-tuning for perceptual enhancement and GRPO-based optimization for reasoning improvement for AVR tasks. Experimental results demonstrate that current MLLMs perform poorly on AVR tasks. Moreover, scaling model parameters, CoT prompting, and inference-time scaling fail to effectively address this challenge. These findings highlight that fine-grained perception constitutes the fundamental bottleneck in AVR, while reasoning enhancements applied based on a perception-grounded model can further improve performance.
In short, our work not only offers a systematic resource for both evaluation and training, but also provides a practical framework for addressing the AVR challenge.
Limitation: Due to limitations in time and human labor, the scale of the resources (e.g., rule design, icon library) used for data synthesis is restricted, which somewhat limits the diversity and richness of the synthetized data. In future work, we plan to expand the resource pool to a larger scale with greater complexity, further enriching the synthesized riddles.
APPENDIX
Appendix A Construction Details of VisuRiddles Benchmark
Collection. We manually collect a diverse set of visual reasoning problems from Chinese National Civil Service Examination, which form the five core categories of the benchmark. RAVEN and Sudoku samples are generated following the method [^1] https://github.com/WellyZhang/RAVEN and the protocols [^2] https://gitee.com/mxx11/sudoku.
Appendix B Details of VisuRiddles Synthesizer
B.1 Configuration of the VisuRiddles Synthesizer
Table 5: Configuration details of the VisuRiddles Synthesizer across seven reasoning categories.
Rule Sub-rule Layout Template Icon
Numerical Line, Curve, Angle, Cart, Space, Parts 1 1 5k
Stylistic AND, OR, XOR, XNOR 3 5 5k
Attribute Element, Group 2 1 5k
Positional Translate, Rotate, Flip 4 7 5k
Spatial Unfolding, Three-View, 3D-Reconstruction 7 13 5k
View-Consistency, Multiple-Views
Sudoku Sudoku 1 5 9
RAVEN Raven 1 7 10k
B.2 Synthesis Pipeline of Riddles Instances
The synthesis pipeline for generating riddle instances with structured perceptual descriptions consists of two main stages: Instance Construction and API Labeling.
Instance Construction
Instance synthesis is responsible for generating the visual instances and corresponding perceptual descriptions. The detailed steps are as follows:
1. Rule Selection:
The synthesis process begins with rule selection. We first extract representative reasoning patterns from real-world standardized tests, such as civil service visual logic exams. These high-level rules are categorized into positional, stylistic, numerical,attribute,spatial,raven and sudoku types, and further decomposed into finer-grained sub-rules, each representing a distinct visual transformation or logical operation (e.g., “simultaneous rotation and revolution”, or “black-white XOR operation”). This decomposition enables modular and interpretable synthesis, allowing each instance to target a specific reasoning skill.
2. Element Configuration:
Once a sub-rule is selected, we proceed to define the visual components in a structured manner. This includes: (i) choosing a layout template (e.g., $3\times 3$ matrix, progression row, etc.), (ii) selecting or auto-generating icon elements with specific visual properties (e.g., shape, color, direction), and (iii) assigning a background grid. Elements are configured in a rule-consistent way: for example, a “positional” sub-rule may require translation across rows, while a “stylistic” sub-rule may apply logical operations to shape fill. Each configuration ensures the pattern remains visually interpretable and solvable.
3. Rendering and Annotating:
API Labeling
Due to the poor performance of existing commercial models in abstract visual reasoning (AVR) tasks, especially when processing images, these models often suffer from significant hallucination issues and fail to generate correct reasoning annotations. Thus, directly feeding images into the model does not produce accurate results. To overcome this, we introduce a new step: first, we replace the images with corresponding perceptual descriptions and input these descriptions into a large model. This allows the model to generate CoT based on the perceptual descriptions. The detailed steps are as follows:
1. CoT Generation:
Perceptual descriptions are first extracted from the synthesized visual instances to supply fine-grained visual information. These descriptions are then provided to an advanced commercial model (Gemini 2.5-Flash-Think [^3] https://deepmind.google/models/gemini/flash/ in this work), which generates a Chain-of-Thought (CoT) reasoning process and produces a corresponding answer.
2. Validation:
The generated answer is compared to the correct answer of the synthesized riddle. If the model’s output does not align with the correct answer, the data point is discarded, ensuring that only valid and accurate samples are included in the final dataset.
B.3 Examples of Synthesized Riddles Instances
Figure 6: Examples of synthesized riddle instances in Numerical, Positional, and Attribute
Figure 7: Examples of synthesized riddle instances in Spatial, Stylistic, Sudoku, and RAVEN
Appendix C Details of Evaluated Models
We evaluate seventeen MLLMs, including eleven open-source models and six advanced proprietary models, as shown in Table  6. Human reference answers are obtained from the Fenbi Educational Platform, a widely used source for civil service exam preparation. Notably, GPT-o1, Claude-3-7-Sonnet-Thinking, and Gemini-2.5-Pro are equipped with specialized “thinking” modes aimed at enhancing multi-step reasoning. To further explore reasoning capabilities, CoT prompting is applied to GPT-4o, Qwen2.5-VL-72B, and InternVL2.5-78B-MPO, selected for their reasoning capacity, model scale, or lack of built-in “thinking” mode. Inference for all models is performed using the vLLM framework with unified decoding configurations.
Table 6: Overview of evaluated MLLMs.
Model Name Params Think CoT Provider Version
Open-Source MLLMs
Qwen2.5-VL-7B [68] 7B Alibaba -
Qwen2.5-VL-32B 32B Alibaba -
Qwen2.5-VL-72B 72B Alibaba -
Qwen2.5-VL-72B(CoT) 72B Alibaba -
Qwen3-VL-235B-A22B-Instruct 235B Alibaba -
Qwen3-VL-235B-A22B-Thinking 235B Alibaba -
InternVL2.5-8B [5] 8B Shanghai AI Lab -
InternVL2.5-MPO-8B [69] 8B Shanghai AI Lab -
InternVL2.5-38B 38B Shanghai AI Lab -
InternVL2.5-38B-MPO 38B Shanghai AI Lab -
InternVL2.5-78B 78B Shanghai AI Lab -
InternVL2.5-78B-MPO 78B Shanghai AI Lab -
InternVL2.5-78B-MPO(CoT) 78B Shanghai AI Lab -
DeepSeek-VL2 [67] 32B DeepSeek -
MiniCPM-V 2.6 [7] 8B ModelBest -
Dots.vlm1 671B Rednote hi lab -
Commercial Products
GPT-4o - OpenAI 2024.11.20
GPT-4o(CoT) - OpenAI 2024.11.20
o3 - OpenAI 2025.04.16
Claude-3.7-Sonnet - Anthropic 2025.02.19
Claude-3.7-Sonnet-Thinking - Anthropic 2025.02.19
Gemini-2.0-Flash-Thinking - Google DeepMind 2025.01.21
Gemini-2.5-Flash-Thinking - Google DeepMind 2025.06.21
Gemini-2.5-Pro - Google DeepMind 2025.03.25
GPT-5 - OpenAI 2025.08.07
Appendix D Experimental result on VisuLogic
Table 7: Experimental result on VisuLogic Benchmark. The best results are marked bold and the second results are underlined.
Models Param. Quantity Spatiality Position Attribute Style Other Overall
Human - 45.3 52.7 71.1 50.0 47.5 44.2 51.4
[…]
Deepseek-R1 670B 27.7 23.5 24.0 27.8 23.0 35.0 26.6
Qwen2.5-72B-Instruct 72B 30.2 24.4 27.5 26.5 26.8 30.8 28.0
Claude-3.7-Sonnet (20250219) - 26.6 22.5 25.0 28.0 25.6 30.6 25.9
Doubao-1.5-Pro-32k (20250115) - 30.0 22.5 25.0 25.6 30.0 24.1 26.6
Close Source MLLMs
GPT-4o-mini (20240718) - 27.2 23.4 23.5 18.3 31.1 16.7 24.3
GPT-4o (20240806) - 28.6 24.7 27.2 26.8 20.0 25.9 26.3
Kimi-latest - 24.9 29.4 26.5 28.0 16.7 26.9 25.9
Doubao-1.5-Vision-Pro-32k (20250115) - 28.1 23.8 29.1 25.1 32.1 35.0 28.1
Gemini-2.0-Pro (20250205) - 29.7 24.2 27.9 30.5 22.2 33.3 28.0
Claude-3.7-Sonnet (20250219) - 22.7 27.3 27.9 28.0 22.2 22.2 24.8
Open Source MLLMs
LLaVA-v1.5 7B 26.1 24.2 23.5 17.1 31.1 22.2 24.6
LLaVA-OneVision (SI) 7B 22.4 27.3 33.1 23.2 25.6 22.2 25.3
Ovis2 8B 26.1 23.8 27.2 28.0 25.6 24.1 25.6
Qwen2.5-VL-7B-Instruct 7B 27.6 20.9 25.2 23.2 37.8 25.0 26.0
Qwen2.5VL-72B-Instruct 72B 25.2 23.8 27.2 25.6 25.6 34.3 26.2
InternVL2.5-38B 8B 24.4 26.4 27.2 23.2 25.6 26.9 25.5
InternVL2.5-78B 78B 26.6 26.0 26.5 26.8 31.1 30.6 27.3
InternVL3-38B 38B 28.7 27.6 26.1 21.4 23.9 28.5 27.1
InternVL3-78B 78B 27.7 26.1 31.6 26.3 21.3 32.3 27.7
Qwen2.5-VL-7B-Instruct-SFT (VisuLogic) 7B 24.4 26.4 27.2 23.2 25.6 26.9 25.5
Qwen2.5-VL-7B-Instruct-RL (VisuLogic) 7B 26.6 33.8 29.4 23.2 18.9 29.6 28.0
InternVL2.5-38B-SFT (VisuLogic) 38B 30.6 29.4 20.6 25.6 30.0 25.0 27.9
InternVL2.5-38B-RL (VisuLogic) 38B 31.2 31.2 26.5 30.5 30.0 38.9 31.1
Ours
PAVR 7B 28.9 30.3 31.6 34.1 30.0 33.3 30.6